OpenStudy (anonymous):

Look at the formula for the volume of a cylinder. V = r2h Which is a solution for h?

OpenStudy (shane_b):

The equation for the volume of a cylinder is \[V=\pi r^2 h\]If you want to isolate h, you can divide both sides by pi r^2 to get:\[\frac{V}{\pi r^2}=h\]

OpenStudy (anonymous):

is that the A?

OpenStudy (shane_b):

I'm not sure what you mean....I see no "A"

OpenStudy (anonymous):

answer

OpenStudy (shane_b):

Oh, yes. That's the equation you're looking for if the question is just asking you to solve for h.

OpenStudy (anonymous):

thnks!